# Cross-modal Full-mode Fine-grained Alignment for Text-to-Image Person Retrieval

### v1

@@ 3.3. Explicit Fine-grained Alignment
To fully exploit fine-grained information, it is crucial to bridge the inherent modality differences between visual and textual modalities. Although most existing fine-grained alignment methods based on attention mechanisms have achieved certain results by implicitly aligning local features of images and texts, they cannot directly observe whether the corresponding local features of image and text are effectively aligned. We introduce an explicit cross-modal aggregation method based on the sparse similarity matrix between the image and text local features. To further reduce redundant information and minimize memory and time costs during fine-grained alignment, we use hard coding to align the aggregated language-grouped vision embeddings with both image and text embeddings, as shown in Figure 2(c).
Sparse Similarity Matrix Aggregation. Some methods (Mukhoti et al., 2023; Yao et al., 2022) use loss functions based on fine-grained representations for cross-modal local alignment, which compute the similarity between all image patch embeddings and text token embeddings. Obviously, this approach is computationally and memory-intensive, making it difficult to scale to large batches. Therefore, some methods attempt to sparsify the full-size similarity matrix and only consider the retrieved embeddings that are similar to the query embedding. Most sparse processing methods are based on softmax, but the softmax approach tends to produce a low-entropy distribution from the perspective of gradient flow (Hoffmann et al., 2024). Therefore, we implement sparse similarity matrix aggregation using max-min normalization.
An image $I$ and its corresponding text $T$ are fed into the visual encoder and textual encoder, respectively, as described in Section 3.1. As illustrated in Figure 3(a), the similarity between all image patches and text tokens is obtained by computing the inner product of the last hidden states $\{f_{i}^{t}\}_{i=1}^{L}$ and $\{f_{i}^{v}\}_{i=1}^{N}$ of the text transformer and the vision transformer. $s_{i,j}=f_{i}^{t}\cdot f_{j}^{v}$ represents the similarity between the text token $f_{i}^{t}$ and the image patch $f_{j}^{v}$, where $\cdot$ denotes the inner product. To obtain the aggregation weight, for each token $i$, we first normalize $s_{i,j}$ to [0,1] using the following min-max normalization formula:
@@
where $\sigma$ is the sparsity threshold. $\sigma$ is set to $1/N$, where $N$ is the number of image patches. This ensures that each token has a minimum of one corresponding image patch for alignment. We compute the aggregation weights by:
@@
where $M$ is the number of image patches retained with high similarity to the token $i$, and $agg_{i,j}$ represents the weight of patch $j$ for aggregating the language-grouped vision embedding corresponding to token $i$. This explicit aggregation strategy ensures a comprehensive interaction between token $i$ and its corresponding patch $j$ during local alignment. In particular, the aggregation weight $agg_{i,j}$ effectively captures the semantic relevance between token $i$ and patch $j$, thereby facilitating precise alignment.
Then, we obtain the corresponding language-grouped vision embedding $e_{i}$ as follows:
@@
where $N$ denotes the number of image patches. The resulting set of language-grouped vision embedding ${e}_{i}$ has the same length $L$ as the text token $f_{i}^{t}$.
Hard Coding Alignment. We calculate the similarity between the language-grouped vision embeddings (referred to as joint embeddings) $\{e_{i}\}_{i=1}^{L}$ and their corresponding original text embeddings $\{f_{i}^{t}\}_{i=1}^{L}$ as well as image embeddings $\{f_{i}^{v}\}_{i=1}^{N}$, respectively. To reduce both computational and memory costs, we adopt a hard coding similarity computation between the joint embeddings and their corresponding text and image embeddings. For simplicity, we only present the calculation between the joint embeddings and the text embeddings, while the remaining computations follow a similar and symmetric approach.
(a) Sparse similarity matrix aggregation.
(b) Hard coding alignment.
Figure 3. Illustration of the EFA module. (a) EFA sparifies the similarity matrix between token embeddings and patch embeddings, and obtains the language-grouped vision embeddings by aggregating the sparse similarity with its corresponding patch. (b) EFA obtains hard similarity through hard coding and LSE pooling, where the calculated hard similarity reflects the relationship between the language-grouped vision embeddings and their original token or patch embeddings.
@@
Then, we utilize the LSE pooling (Lee et al., 2018) to compute the hard similarity between text $T$ and its corresponding joint embedding $E$ by:
@@
where $\lambda$ is a factor that determines how much to magnify the importance of the most relevant pairs of text embedding and its corresponding joint embedding.
Given a mini-batch of $B$ texts and its corresponding joint embeddings, we obtain the hard coding similarity matrix Hard_SHard\_S according to Eq. (9) and Eq. (10), as illustrated in Figure 3(b). We calculate the EFA loss from the text to its joint embedding, adapted from the triplet ranking loss (Schroff et al., 2015):
@@
where $\tau_{2}$ is a temperature hyperparameter that controls the distribution of the loss, and marginmargin is a distance hyperparameter representing the minimum distance difference between positive and negative samples.
Similarly, the EFA loss from the joint embedding to its original text can be computed following Eq. (11), and we can calculate the EFA loss between image and its joint embedding according to Eq. (9), Eq. (10) and Eq. (11). Then, we obtain the full EFA loss by:
@@
Theoretical analysis of hard coding: Given visual embeddings of an image $\boldsymbol{V}\in\mathbb{R}^{N\times d}$ and its joint embeddings $\boldsymbol{E}\in\mathbb{R}^{L\times d}$, where $N$ is the number of image patches and $L$ is the length of the joint embeddings, both the hard coding method and the soft coding method calculate the similarity matrix $A$ to obtain the final similarity between $\boldsymbol{V}$ and $\boldsymbol{E}$, which requires the same time complexity $\mathcal{O}(NLd)$. When processing the similarity matrix $A$, our hard coding method calculates the maximum similarity for each row, which has a time complexity of $\mathcal{O}(N)$, as described in Eq. (9) and Eq. (10). However, the soft coding method processes all similarities, resulting in a time complexity of $\mathcal{O}(NL)$. Moreover, the spatial complexity of the hard coding method is $\mathcal{O}(NL)$, which is significantly lower than the spatial complexity of the soft coding method, $\mathcal{O}(Ld)$. This is based on the fact that $N\ll d$. Therefore, our hard coding method can effectively reduce the cost of training time and memory.
@@ 3.4. Training Objective
As mentioned, the primary goal of FMFA is to enhance the global and cross-modal local alignment of image-text representations within the joint embedding space. To achieve it, the most commonly used ID loss (Zheng et al., 2020) and IRR loss (Jiang and Ye, 2023), together with the proposed EFA loss and the A-SDM loss, are jointly employed to optimize FMFA. The ID loss directly classifies the global features extracted from each image and text according to their identities, thereby enhancing the global alignment of the model. The IRR loss, based on the Masked Language Modeling (MLM) task (Taylor, 1953), leverages an attention mechanism for implicit cross-modal interaction to obtain a joint embedding, and then predicts the [MASK] text token to enhance the local alignment of the model.
FMFA is trained in an end-to-end manner and the overall training objective is defined as:
@@ 4.1. Datasets and Settings
In this section, we present the datasets, evaluation metrics and implementation details in our experiments.
Datasets. We use the CHUK-PEDES (Li et al., 2017), ICFG-PEDES (Ding et al., 2021), and RSTPReid (Zhu et al., 2021) datasets to train and evaluate our FMFA. We follow the data partitions used in IRRA (Jiang and Ye, 2023) to split each dataset into training, validation and test sets, where the ICFG-PEDES dataset only has training and validation sets.
  • CUHK-PEDES is the first benchmark specifically designed for text-to-image person retrieval. It comprises 40,206 images and 80,412 textual descriptions corresponding to 13,003 identities. According to the official split, the training set includes 11,003 identities, with 34,054 images and 68,108 descriptions. The validation and test sets contain 3,078 and 3,074 images, as well as 6,158 and 6,156 textual descriptions, respectively, each covering 1,000 distinct identities.
  • ICFG-PEDES consists of 54,522 images corresponding to 4,102 identities, with each image paired with a single textual description. The dataset is split into a training set and a test set: the training set includes 34,674 image-text pairs from 3,102 identities, while the test set comprises 19,848 pairs covering the remaining 1,000 identities.
  • RSTPReid comprises 20,505 images of 4,101 identities captured by 15 different cameras. Each identity is associated with 5 images taken from different viewpoints, and each image is annotated with 2 textual descriptions. According to the official split, the training, validation, and test sets contain 3,701, 200, and 200 identities, respectively.
Evaluation Metrics. We mainly employ the popular Rank-K metrics (K=1, 5, 10) to measure the retrieval performance. In addition to Rank-K, we also adopt the mean Average Precision (mAP) as an auxiliary retrieval metric to further evaluate performance. Higher evaluation metrics indicate better performance of the model.
@@ 4.2. Comparison with State-of-the-Art Methods
In this section, we present a comparison with state-of-the-art methods (e.g., NAM (Tan et al., 2024) and HAM (Jiang et al., 2025)) on three public benchmark datasets. Based on the backbone type, we categorize the baselines into two groups, as shown in Table 2, Table 3, and Table 4: those employing VL-Backbones w/o ReID-domain pre-training and with ReID-domain pre-training. Furthermore, according to whether local features are utilized during inference, the baselines are further divided into local matching methods and global matching methods (denoted as “L” and “G” in the “Type” column, respectively). Note that the baseline model in Table 2, Table 3 and Table 4 is denoted as IRRAR, which represents the performance of our reimplementation of the IRRA model. CLIP refers to the CLIP-ViT-B/16 model fine-tuned with the InfoNCE loss (Oord et al., 2018).
Table 2. Performance comparisons with state-of-the-art methods on the CUHK-PEDES dataset. “G”, “L” and “P” in the “Type” column stand for global-matching method, local-matching method and pre-trained model with ReID-domain respectively. “Image Enc.” and “Text Enc.” mean the backbone of image encoder and text encoder respectively. “IRRAR” means the model that we reproduce.
@@
G LGUR (Shao et al., 2022) MM22 ResNet50 BERT 65.25 83.12 89.00 -
IVT (Shu et al., 2022) ECCV22 ViT-B/16 BERT 65.59 83.11 89.21 -
CLIP (Radford et al., 2021) ICML21 CLIP-ViT CLIP-Xformer 68.19 86.47 91.47 61.12
DM-Adapter (Liu et al., 2025) AAAI25 CLIP-ViT CLIP-Xformer 72.17 88.74 92.85 64.33
IRRAR (Jiang and Ye, 2023) CVPR23 CLIP-ViT CLIP-Xformer 73.45 89.38 93.69 66.13
TBPS-CLIP (Cao et al., 2024) AAAI24 CLIP-ViT CLIP-Xformer 73.54 88.19 92.35 65.38
@@
L ACSA (Ji et al., 2022) TMM22 Swin-B BERT 63.56 81.49 87.70 -
Han et al. (Han et al., 2021) arXiv21 CLIP-RN101 CLIP-Xformer 64.08 81.73 88.19 60.08
PLOT (Park et al., 2024) ECCV24 CLIP-ViT CLIP-Xformer 75.28 90.42 94.12 -
RaSa (Bai et al., 2023b) IJCAI23 CLIP-ViT BERT-base 76.51 90.29 94.25 69.38
PTMI (Lu et al., 2025) TIFS25 CLIP-ViT CLIP-Xformer 76.02 89.93 94.14 70.85
APTM (Yang et al., 2023) MM23 Swin-B BERT-base 76.53 90.04 94.15 66.91
SCVD (Wei et al., 2024) TCSVT24 CLIP-RN50 CLIP-Xformer 76.72 90.38 94.89 -
@@
P+G UniPT (Shao et al., 2023) + IRRA (Jiang and Ye, 2023) ICCV23 CLIP-ViT CLIP-Xformer 74.37 89.51 93.97 66.60
PLIP (Zuo et al., 2024) + IRRA (Jiang and Ye, 2023) NeurIPS24 CLIP-ViT CLIP-Xformer 74.25 89.49 93.68 66.52
NAM (Tan et al., 2024) + IRRAR CVPR24 CLIP-ViT CLIP-Xformer 76.67 91.11 94.60 68.42
NAM (Tan et al., 2024) + FMFA (ours) CLIP-ViT CLIP-Xformer 77.23 91.33 94.75 68.53
HAM (Jiang et al., 2025) + IRRAR CVPR25 CLIP-ViT CLIP-Xformer 77.32 91.20 94.95 68.87
HAM (Jiang et al., 2025) + FMFA (ours) CLIP-ViT CLIP-Xformer 77.46 91.36 95.01 68.89
Performance Comparisons on CUHK-PEDES. We evaluate FMFA on the widely used CUHK-PEDES benchmark, as shown in Table 2. When using the VL-Backbones without ReID-domain pre-training, FMFA outperforms all state-of-the-art global matching methods with 74.16% Rank-1 accuracy and 66.66% mAP, and further exceeds IRRA by 0.74% and 0.41% in Rank-5 and Rank-10, respectively. When adopting the VL-Backbones with ReID-domain pre-training, FMFA maintains its superiority, and achieves Rank-5 accuracy exceeding 95% with the HAM-based backbone. Notably, FMFA with NAM-based backbone attains 91.33% in Rank-5, outperforming IRRA with the HAM-based backbone by 0.13%.
Table 3. Performance comparisons with state-of-the-art methods on the RSTPReid dataset.
@@
G DSSL (Zhu et al., 2021) 39.05 62.60 73.95 -
IVT (Shu et al., 2022) 46.70 70.00 78.80 -
CLIP (Radford et al., 2021) 54.05 80.70 88.00 43.41
IRRAR (Jiang and Ye, 2023) 59.50 81.80 88.85 47.44
DM-Adapter (Liu et al., 2025) 60.00 82.10 87.90 47.37
@@
L ACSA (Lu et al., 2025) 48.40 71.85 81.45 -
CFine (Yan et al., 2023a) 50.55 72.50 81.60 -
PLOT (Park et al., 2024) 61.80 82.85 89.45 -
RaSa (Bai et al., 2023b) 66.90 86.50 91.35 52.31
APTM (Yang et al., 2023) 67.50 85.70 91.45 52.56
@@
P+G UniPT (Shao et al., 2023) + IRRA (Jiang and Ye, 2023) 62.20 83.30 89.75 48.33
PLIP (Zuo et al., 2024) + IRRA (Jiang and Ye, 2023) 64.35 83.75 91.00 50.93
NAM (Tan et al., 2024) + IRRAR 68.25 86.75 92.30 52.92
NAM (Tan et al., 2024) + FMFA (ours) 68.70 87.05 92.35 53.14
HAM (Jiang et al., 2025) + IRRAR 71.35 87.60 93.05 55.40
HAM (Jiang et al., 2025) + FMFA (ours) 71.80 88.05 93.15 55.72
Performance Comparisons on RSTPReid. We evaluate FMFA on the newest benchmark, RSTPReid, as shown in Table 3. Using the VL-Backbones without ReID-domain pre-training, FMFA achieves competitive performance, attaining 61.05% Rank-1, 83.85% Rank-5, 89.80% Rank-10, and 48.22% mAP, respectively, outperforming IRRA by 1.55% in Rank-1 and 2.05% in Rank-5. When adopting the VL-Backbones with ReID-domain pre-training, our method achieves further gains, exceeding IRRA by 0.45% in Rank-1 with both the NAM-based and HAM-based backbones. Notably, FMFA achieves Rank-5 accuracy higher than 88% with the HAM-based backbone.
Table 4. Performance comparisons with state-of-the-art methods on the ICFG-PEDES dataset.
@@
G Dual Path (Zheng et al., 2020) 38.99 59.44 68.41 -
IVT (Shu et al., 2022) 56.04 73.60 80.22 -
CLIP (Radford et al., 2021) 56.74 75.72 82.26 31.84
DM-Adapter (Liu et al., 2025) 62.64 79.53 85.32 36.50
IRRAR (Jiang and Ye, 2023) 63.48 80.16 85.78 38.20
@@
L SSAN (Ding et al., 2021) 54.23 72.63 79.53 -
ISANet (Yan et al., 2023b) 57.73 75.42 81.72 -
CFine (Yan et al., 2023a) 60.83 76.55 82.42 -
RaSa (Bai et al., 2023b) 65.28 80.40 85.12 41.29
PLOT (Park et al., 2024) 65.76 81.39 86.73

### v2

@@ 3.3. Explicit Fine-grained Alignment
To effectively leverage fine-grained information, it is necessary to narrow the underlying disparity between visual and textual modalities. Although many attention-based fine-grained alignment approaches have shown effectiveness by implicitly associating local regions in images with textual fragments, they provide no direct means to verify whether these localized correspondences are accurately aligned. We propose an explicit cross-modal aggregation approach that leverages the sparse similarity matrix between the local image and text features. To further reduce redundant information and minimize memory and time costs during fine-grained alignment, we use hard coding to align the aggregated language-grouped vision embeddings with both image and text embeddings, as shown in Figure 2(c).
(a) Sparse similarity matrix aggregation.
(b) Hard coding alignment.
Figure 3. Illustration of the EFA module. (a) EFA imposes a sparse structure on the similarity matrix relating token embeddings to patch embeddings, and obtains the language-grouped vision embeddings by aggregating the sparse similarity with its corresponding patch. (b) EFA obtains hard similarity through hard coding and LSE pooling, where the calculated hard similarity reflects the relationship between the language-grouped vision embeddings and their original token or patch embeddings.
Sparse Similarity Matrix Aggregation. Some methods (mukhoti2023open; yao2021filip) incur substantial computational and memory overhead, as they evaluate pairwise relationships between every image patch and every text token, which limits scalability to large batch sizes. Therefore, we apply a sparsification strategy to reduce the full pairwise similarity computation. While softmax is commonly used for such sparse processing, it tends to produce low-entropy similarity distributions that impede effective gradient flow (hoffmann2023eureka). Thus, we further adopt a max-min normalization scheme to achieve a more stable and expressive sparse similarity aggregation.
An image $I$ and its corresponding text $T$ are encoded through the visual and textual encoders, respectively. As presented in Figure 3(a), the similarity between image patches and text tokens is computed via the inner product of the last hidden states $\{f_{i}^{t}\}_{i=1}^{L}$ of the text transformer and $\{f_{i}^{v}\}_{i=1}^{N}$ of the vision transformer. $s_{i,j}=f_{i}^{t}\cdot f_{j}^{v}$ measures the similarity between the text token $f_{i}^{t}$ and the image patch $f_{j}^{v}$, where $\cdot$ denotes the inner product. To obtain the aggregation weight, each token $i$ is first scaled to the range [0,1] through the following min-max normalization:
@@
where $\sigma$ is the sparsity threshold. $\sigma$ is assigned the value $1/N$, where $N$ corresponds to the total count of patches in the image. This ensures that each token has a minimum of one corresponding image patch for alignment. We compute the aggregation weights by:
@@
where $M$ is the number of image patches retained with high similarity to the token $i$, and $agg_{i,j}$ quantifies the influence of patch $j$ in forming the language-grouped vision embedding (referred to as joint embedding) associated with token $i$. This explicit aggregation strategy ensures a comprehensive interaction between token $i$ and its corresponding patch $j$ during local alignment. In particular, the aggregation weight $agg_{i,j}$ effectively captures the semantic relevance between token $i$ and patch $j$, thereby facilitating precise alignment.
Next, we derive the corresponding joint embedding $e_{i}$ as:
@@
where $N$ is the count of image patches. The resulting set of joint embedding ${e}_{i}$ has the same length $L$ as the text token $f_{i}^{t}$.
Hard Coding Alignment. We calculate the similarity between the joint embeddings $\{e_{i}\}_{i=1}^{L}$ and their corresponding original text embeddings $\{f_{i}^{t}\}_{i=1}^{L}$ as well as image embeddings $\{f_{i}^{v}\}_{i=1}^{N}$, respectively. To reduce both computational and memory costs, we adopt a hard coding similarity computation between the joint embeddings and their corresponding text and image embeddings, and the theoretical analysis of the hard coding is provided in Appendix A. For simplicity, we only present the calculation between the joint embeddings and the text embeddings, while the remaining computations follow a similar and symmetric approach.
@@
Then, we utilize the LSE pooling (lee2018stacked) to compute the hard similarity between text $T$ and its corresponding joint embedding $E$ by:
@@
where $\lambda$ controls the degree to which the most relevant text embeddings and their corresponding joint embeddings are emphasized.
Given a batch containing $B$ text embeddings along with their associated joint embeddings, we compute the hard coding similarity matrix Hard_SHard\_S following Eq. (9) and Eq. (10), as illustrated in Figure 3(b). We calculate the EFA loss from the text to its joint embedding, adapted from the triplet ranking loss (schroff2015facenet):
@@
where $\tau_{2}$ is a scaling factor adjusting the spread of the loss, and marginmargin is a distance hyperparameter defining the minimal gap separating positive and negative pairs.
Similarly, the EFA loss from the joint embedding to its original text can be computed following Eq. (11), and we can calculate the EFA loss between image and its joint embedding through Eq. (9), Eq. (10) and Eq. (11). Then, we obtain a full EFA loss by:
@@ 3.4. Training Objective
As mentioned, FMFA aims to improve both the global and local cross-modal alignment of image-text features within the shared embedding space. To realize this goal, the widely adopted ID loss (zheng2020dual) and IRR loss (jiang2023cross), together with the proposed EFA and A-SDM loss, are jointly utilized to train FMFA. The ID loss directly classifies the global features obtained from both the image and the text according to their identities, thereby enhancing the global alignment of the model. The IRR loss, based on the Masked Language Modeling (MLM) task (taylor1953cloze), leverages an attention mechanism for implicit cross-modal interaction to obtain a joint embedding, and then predicts the [MASK] text token to enhance the local alignment of the model.
FMFA is trained end-to-end, with the complete training objective formulated as:
@@ 4.1. Datasets and Settings
Datasets. We assess FMFA on three widely used text-based person retrieval datasets, following the data splits introduced in IRRA (jiang2023cross). CUHK-PEDES (li2017person) contains 40,206 images associated with 13,003 identities, where each image is paired with two textual descriptions. Of these identities, 11,003 are designated for training, while the remaining 1,000 identities are allocated separately to validation and test sets. ICFG-PEDES (ding2021semantically) includes 54,522 images belonging to 4,102 individuals, each image linked to a single sentence. The conventional setup utilizes 3,102 identities for training and reserves 1,000 identities for testing. RSTPReid (zhu2021dssl) comprises 20,505 images from 4,101 identities captured across 15 camera views. Every identity corresponds to five images taken from different viewpoints, and each image is annotated with two descriptive captions. The dataset follows a split with 3,701 identities for training and 200 identities each for validation and testing.
Evaluation Metrics. To gauge retrieval quality, we primarily report Rank-K results (K = 1, 5, 10), which measure how often the correct item appears within the top-K predictions. Additionally, mean Average Precision (mAP) is adopted to summarize ranking accuracy over all query outcomes. In both cases, higher metric values correspond to superior model behavior.
@@
Table 2. The hardware configuration of our experimental environment.
Hardware Details
CPU Intel Xeon Gold 6330
GPU NVIDIA RTX A6000
RAM 755 GB DDR4
@@ 4.2. Comparison with State-of-the-Art Methods
In this subsection, we provide a comparison with current state-of-the-art methods (e.g., NAM (tan2024harnessing) and HAM (jiang2025modeling)) on three public benchmark datasets. The methods are grouped into two types according to their underlying network architecture, as listed in Table 3, Table 4, and Table 5: those using VL-Backbones without ReID-domain pre-training and those incorporating ReID-domain pre-training. Furthermore, according to whether local features are utilized during inference, the baselines are further classified into local and global matching methods (denoted as “L”” and “G” in the “Type” column, respectively). It should be noted that the baseline model presented in Table 3, Table 4, and Table 5 is referred to as IRRAR, which represents the performance of our reimplementation of the IRRA model. CLIP means the ViT-B/16 architecture after fine-tuning under the InfoNCE loss (oord2018representation).
Table 3. Comparisons with state-of-the-art methods on the CUHK-PEDES dataset. “G”, “L” and “P” in the “Type” column stand for global-matching method, local-matching method and pre-trained model with ReID-domain respectively. “Image Enc.” and “Text Enc.” mean the backbone of image encoder and text encoder respectively. “IRRAR” means the model that we reproduce.
@@
G LGUR (shao2022learning) MM22 ResNet50 BERT 65.25 83.12 89.00 -
IVT (shu2022see) ECCV22 ViT-B/16 BERT 65.59 83.11 89.21 -
VGSG (he23vgsg) TIP23 ResNet50 Transformer 67.52 84.37 90.26 -
CLIP (radford2021learning) ICML21 CLIP-ViT CLIP-Xformer 68.19 86.47 91.47 61.12
DM-Adapter (liu2025dm) AAAI25 CLIP-ViT CLIP-Xformer 72.17 88.74 92.85 64.33
IRRAR (jiang2023cross) CVPR23 CLIP-ViT CLIP-Xformer 73.45 89.38 93.69 66.13
TBPS-CLIP (cao2024empirical) AAAI24 CLIP-ViT CLIP-Xformer 73.54 88.19 92.35 65.38
@@
L ACSA (ji2022asymmetric) TMM22 Swin-B BERT 63.56 81.49 87.70 -
Han et al. (han2021text) arXiv21 CLIP-RN101 CLIP-Xformer 64.08 81.73 88.19 60.08
PLOT (park2024plot) ECCV24 CLIP-ViT CLIP-Xformer 75.28 90.42 94.12 -
RaSa (bai2023rasa) IJCAI23 CLIP-ViT BERT-base 76.51 90.29 94.25 69.38
PTMI (lu2025prompt) TIFS25 CLIP-ViT CLIP-Xformer 76.02 89.93 94.14 70.85
APTM (yang2023towards) MM23 Swin-B BERT-base 76.53 90.04 94.15 66.91
SCVD (wei2024fine) TCSVT24 CLIP-RN50 CLIP-Xformer 76.72 90.38 94.89 -
@@
P+G UniPT (shao2023unified) + IRRA (jiang2023cross) ICCV23 CLIP-ViT CLIP-Xformer 74.37 89.51 93.97 66.60
PLIP (zuo2024plip) + IRRA (jiang2023cross) NeurIPS24 CLIP-ViT CLIP-Xformer 74.25 89.49 93.68 66.52
NAM (tan2024harnessing) + IRRAR CVPR24 CLIP-ViT CLIP-Xformer 76.67 91.11 94.60 68.42
NAM (tan2024harnessing) + FMFA (ours) CLIP-ViT CLIP-Xformer 77.23 91.33 94.75 68.53
HAM (jiang2025modeling) + IRRAR CVPR25 CLIP-ViT CLIP-Xformer 77.32 91.20 94.95 68.87
HAM (jiang2025modeling) + FMFA (ours) CLIP-ViT CLIP-Xformer 77.46 91.36 95.01 68.89
Table 4. Comparisons with state-of-the-art methods on the RSTPReid dataset.
@@
G DSSL (zhu2021dssl) 39.05 62.60 73.95 -
IVT (shu2022see) 46.70 70.00 78.80 -
CLIP (radford2021learning) 54.05 80.70 88.00 43.41
IRRAR (jiang2023cross) 59.50 81.80 88.85 47.44
DM-Adapter (liu2025dm) 60.00 82.10 87.90 47.37
@@
L ACSA (lu2025prompt) 48.40 71.85 81.45 -
CFine (yan2023clip) 50.55 72.50 81.60 -
PLOT (park2024plot) 61.80 82.85 89.45 -
RaSa (bai2023rasa) 66.90 86.50 91.35 52.31
APTM (yang2023towards) 67.50 85.70 91.45 52.56
@@
P+G UniPT (shao2023unified) + IRRA (jiang2023cross) 62.20 83.30 89.75 48.33
PLIP (zuo2024plip) + IRRA (jiang2023cross) 64.35 83.75 91.00 50.93
NAM (tan2024harnessing) + IRRAR 68.25 86.75 92.30 52.92
NAM (tan2024harnessing) + FMFA (ours) 68.70 87.05 92.35 53.14
HAM (jiang2025modeling) + IRRAR 71.35 87.60 93.05 55.40
HAM (jiang2025modeling) + FMFA (ours) 71.80 88.05 93.15 55.72
Evaluation Results on CUHK-PEDES We measure the performance of FMFA on the CUHK-PEDES dataset, as presented in Table 3. When using the VL-Backbones without ReID-domain pre-training, FMFA achieves superior performance over advanced global matching methods, attaining 74.16% Rank-1 and 66.66% mAP, while surpassing IRRA by 0.74% in Rank-5 and 0.41% in Rank-10. When adopting the VL-Backbones with ReID-domain pre-training, FMFA maintains its superiority, and achieves Rank-5 accuracy exceeding 95% with the HAM-based backbone. Notably, FMFA with NAM-based backbone attains 91.33% in Rank-5, outperforming IRRA with the HAM-based backbone by 0.13%.
Evaluation Results on RSTPReid. We assess FMFA on the latest RSTPReid benchmark, as presented in Table 4. Using the VL-Backbones without ReID-domain pre-training, FMFA achieves competitive performance, attaining 61.05% Rank-1, 83.85% Rank-5, 89.80% Rank-10, and 48.22% mAP, respectively, outperforming IRRA by 1.55% in Rank-1 and 2.05% in Rank-5. When adopting the VL-Backbones with ReID-domain pre-training, our method achieves further gains, exceeding IRRA by 0.45% in Rank-1 with both the NAM-based and HAM-based backbones. Notably, FMFA achieves Rank-5 accuracy higher than 88% with the HAM-based backbone.
Table 5. Comparisons with state-of-the-art methods on the ICFG-PEDES dataset.
@@
G Dual Path (zheng2020dual) 38.99 59.44 68.41 -
IVT (shu2022see) 56.04 73.60 80.22 -
CLIP (radford2021learning) 56.74 75.72 82.26 31.84
VGSG (he23vgsg) 60.34 76.01 82.01 -
DM-Adapter (liu2025dm) 62.64 79.53 85.32 36.50
IRRAR (jiang2023cross) 63.48 80.16 85.78 38.20
@@
L SSAN (ding2021semantically) 54.23 72.63 79.53 -
ISANet (yan2023image) 57.73 75.42 81.72 -
CFine (yan2023clip) 60.83 76.55 82.42 -
RaSa (bai2023rasa) 65.28 80.40 85.12 41.29
PLOT (park2024plot) 65.76 81.39 86.73